OpenStudy (anonymous):

write an equation in point slope form for the perpendicular bisector of the segment with endpoints M(-1,-3)and N (7,1)

OpenStudy (amistre64):

slope of segment here = 4/8 = 1/2 slope of perp needs to equal -2 ..................................................... midpoint of segment = (3,-1) so include that as a point: y = -2(x-3)+1 y = -2x + 7 maybe?

OpenStudy (amistre64):

point slope form...ok (y--1) = -2(x-3)

OpenStudy (amistre64):

so my slope int form was off by 2 :)

OpenStudy (anonymous):

thanks so much! =)